OpenStudy (anonymous):

What is the solution to the equation 36^3c = square root of 6^ (6c – 8) ?

OpenStudy (anonymous):

this your equation?: \[\large 36^{3c}=\sqrt{6^{6c-8}}\]

OpenStudy (anonymous):

\[\large 36^{3c}=\sqrt{6^{6c-8}}\] \[\large (6^2)^{3c}=\sqrt{6^{6c-8}}\] \[\large (6)^{6c}=(6^{6c-8})^{1/2}\] \[\large 6^{6c}=6^{3c-4}\] \[6c=3c-4\] now just solve for c...